The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by venugopalramana(3286)

The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
let base =b
8 cm less than this =b-8 =one equal leg of triangle
perimeter=b+b-8+b-8=3b-16=29
3b=29+16=45
b=45/3=15
equal leg =b-8=15-8=7